# Truck Parking Usage Prediction with Decomposed Graph Neural Networks

### v2

@@ IV-F Regional vs Random Graph Decomposition
IV-G Discussion
@@
The experiment results show that both the regional and random decomposition technique outperform the non-decomposed method, T-GCN. This suggests that the concatenation of hidden vectors plays a crucial role in sharing subgraph features, highlighting the importance of employing graph neural networks in conjunction with pre-clustering graphs for superior performance.
@@
As a potential future research, we propose exploring the concept of decomposing by highway functions, which is more closely related to the behaviors of truck drivers. Graph connectivity also needs to be reconfigured aligned with truck drivers decision making process, which they tend to seek truck parking sites on the same or close highway.
The investigation of the limits of over-smoothing [33] on our models is also important. By incorporating deeper layers and exploring the potential for extending prediction time horizon, we will find our model’s maximum allowable performance limits. Lastly, it is necessary to design an end-to-end framework capable of automatically decomposing graphs and integrating spatio-temporal modules for inductive scenarios. This will enable us to apply the models to previously unseen graphs, such as those outside the scope of participating states in MAASTO while maintaining consistent and robust performance. These future directions hold the potential to further advance the capabilities and applicability of our proposed framework.
@@ V Conclusion
In this paper, we propose the Regional Temporal Graph Neural Network (RegT-GCN), a state-wide truck parking usage prediction framework, leveraging topological structures of truck parking site distributions. The RegT-GCN comprises a Regional Decomposition approach to learn the geographical characteristics efficiently and a spatio-temporal module to capture the temporal dependencies on graphs. Experiments using our original field dataset show that the proposed models outperform other baseline models. Decomposed graphs into regional-oriented subgraphs result in better prediction capability and consistency of inferences. The study also encompasses a demonstration of predicting usages through the empirical observation of non-informative loss, facilitated by the integration of the Gated Recurrent Unit module, which adeptly adapts to value fluctuations. Furthermore, with extensive experiments focused on graph connectivity, this paper validates that incorporating regional relationships of truck parking enhances the comprehension of the structural dynamics of truck parking behavior. In future work, we will explore better graph connectivity based on highway functions in an end-to-end manner.

### v3

@@ IV-F Regional vs Random Graph Decomposition
IV-G Parameters Sensitivity Analysis
Lastly, we examined the model sensitivity by varying the input time step as {1, 3, 6, 9, 12}, the hidden layers dimension as {16, 32, 64, 128, 256, 512}, and the output dimension of RegT-GCN as {16, 32, 64, 128, 256, 512} (Figure 6). When testing the hidden layer dimensions, the output dimensions were fixed at 256, and vice versa.
The first analysis (top) demonstrates that shorter input time steps initially produce lower RMSE, and the results are not always consistent across different output horizons, although longer input sequences are often expected to yield better predictive performance. Interestingly, the RMSE values for all input configurations converge as the output time steps increase. This convergence indicates that, as the model predicts farther into the future, the accumulated prediction uncertainty limits the impact of the initial input length. Similarly, we can assume that increasing the input length yields diminishing results at a certain point when the input length is approximately balanced.
The second analysis (bottom left) indicates that while increasing hidden layer dimensions generally reduces RMSE, there are fluctuations at 256 dimensions. Finally, the third analysis (bottom right) shows that increasing the RegT-GCN output dimensions improves the accuracy, with RMSE decreasing up to 256 dimensions, after which the improvements plateau. While larger output spaces allow richer feature representations, there is a point beyond which additional complexity yields diminishing returns. These results suggest that model architecture must be carefully tuned to balance performance, as overly complex networks may not always improve results.
IV-H Discussion
IV-H1 Computational Efficiency
@@
IV-H2 Graph Connectivity
The experiment results show that both the regional and random decomposition techniques outperform the non-decomposed method, T-GCN. This suggests that the concatenation of hidden vectors plays a crucial role in sharing subgraph features, highlighting the importance of employing graph neural networks in conjunction with pre-clustering graphs for superior performance.
@@ IV-H2 Graph Connectivity
IV-I Limitations and Future Work
Route Planning and Graph Connectivity: Corresponding to the truck driver’s decision-making process, our research is limited to exploring the graph connection along with their route planning. We will explore the concept of decomposition by their route planning network in our future research.
Over-smoothing: The investigation of the limits of over-smoothing [41] on our models is also important. By incorporating deeper layers and exploring the potential for extending the prediction time horizon, the proposed model’s maximum allowable performance limits should be investigated in the future.
Model Composability: It is necessary to design an end-to-end framework capable of automatically decomposing graphs and integrating spatio-temporal modules for inductive scenarios. This will enable us to apply the models to previously unseen graphs, such as those outside the scope of participating states in MAASTO, while maintaining consistent and robust performance.
@@ V Conclusion
In this paper, we propose the Regional Temporal Graph Convolutional Network (RegT-GCN), a state-wide truck parking usage prediction framework, leveraging topological structures of truck parking site distributions. The RegT-GCN comprises a Regional Decomposition approach to learn the geographical characteristics efficiently and a spatio-temporal module to capture the temporal dependencies on graphs. Experiments using our original field dataset show that the proposed models outperform other baseline models. Decomposed graphs into regional-oriented subgraphs result in better prediction capability and consistency of inferences. The study also encompasses a demonstration of predicting usages through the empirical observation of non-informative loss, facilitated by the integration of the Gated Recurrent Unit module, which adeptly adapts to value fluctuations. Furthermore, with extensive experiments focused on graph connectivity, this paper validates that incorporating regional relationships of truck parking enhances the comprehension of the structural dynamics of truck parking behavior.